# Are Existing Road Design Guidelines Suitable for Autonomous Vehicles?

### v1

@@ Abstract
The emergence of Autonomous Vehicles (AVs) has spurred research into testing the resilience of their perception systems, i.e. to ensure they are not susceptible to making critical misjudgements. It is important that they are tested not only with respect to other vehicles on the road, but also those objects placed on the roadside. Trash bins, billboards, and greenery are all examples of such objects, typically placed according to guidelines that were developed for the human visual system, and which may not align perfectly with the needs of AVs. Existing tests, however, usually focus on adversarial objects with conspicuous shapes/patches, that are ultimately unrealistic given their unnatural appearances and the need for white box knowledge. In this work, we introduce a black box attack on the perception systems of AVs, in which the objective is to create realistic adversarial scenarios (i.e. satisfying road design guidelines) by manipulating the positions of common roadside objects, and without resorting to ‘unnatural’ adversarial patches. In particular, we propose $\mathtt{TrashFuzz}$, a fuzzing algorithm to find scenarios in which the placement of these objects leads to substantial misperceptions by the AV—such as mistaking a traffic light’s colour—with overall the goal of causing it to violate traffic laws. To ensure the realism of these scenarios, they must satisfy several rules encoding regulatory guidelines about the placement of objects on public streets. We implemented and evaluated these attacks for the Apollo, finding that $\mathtt{TrashFuzz}$ induced it into violating 15 out of 24 different traffic laws.
@@
Autonomous Driving System, Autonomous Vehicle, Testing, Security, Fuzzing.
1 Introduction
Autonomous Vehicles (AVs) are currently undergoing rapid and promising development. Notably, several Level-4 AVs—which do not require driver intervention—have been successfully deployed in real-world traffic scenarios [1], including Google Waymo [2], Baidu Apollo [3], and TuSimple [4]. This success has spurred research into testing the resilience of their perception systems, i.e. to ensure they are not susceptible to making critical misjudgements. It is important that they are tested not only with respect to vehicles and obstacles on the road, but also those objects placed to its side, which have the potential to affect the overall perception. Trash bins, billboards, and greenery are all examples of such objects, typically placed according to guidelines (e.g. [5, 6, 7, 8, 9, 10]) that were developed for the human visual system, and may not align perfectly with the requirements of AVs.
Existing tests and attacks, however, usually focus on adversarial objects with conspicuous shapes or patches that have the potential to be noticed by human observers (e.g. [11, 12, 13, 14, 15, 16]). For instance, the adversarial traffic cone in MSF-ADV [11] significantly deviates from the appearance of a regular traffic cone, rendering it as relatively easy for humans to identify as suspicious. Furthermore, the objectives of current attacks tend to be limited to goals such as “ignoring the object” [11] or “misjudging the position of the object” [13, 14, 15, 16]. While these objectives could occasionally result in safety-critical situations, such as a high-speed collision with a traffic cone, the consequences of these misperceptions may be relatively benign with respect to the safety of the overall journey. Finally, these attacks generally require white box knowledge of the underlying models in the perception system, which is difficult to obtain in practice. Ultimately, these limitations make it challenging for existing attacks to be applied successfully on real roads.
These observations led us to consider two questions. First, to what extent can AV perception systems be deceived by objects that follow conventional road design guidelines and appear to humans as ‘natural’? In other words, is there a risk that certain placements of everyday (patch-free) roadside objects could accidentally (or intentionally) trigger a misperception by the AV? Second, if such a misperception occurs, is there a risk that it could lead to serious safety issues on the road, e.g. a collision, or even the violation of traffic laws [17]? If such scenarios are feasible—even if likely to be rare—then it is important to be able to demonstrate their presence in a testing tool before the AV encounters them on real roads.
@@ 1 Introduction
Figure 1: $\mathtt{TrashFuzz}$ workflow: the fuzzer mutates the placement of roadside objects (e.g. trash bins, trees) so as to find ‘natural’ scenarios that induce the AV to violate traffic laws
To answer these questions, we developed a black box approach for systematically fuzzing the perception systems of AVs and uncovering scenarios in which their Autonomous Driving Systems (ADSs) are deceived into making dangerous manoeuvres. In particular, we generate scenarios consisting of ‘naturally’ placed, shaped, and patch-free roadside objects such as trash bins, benches, trees, and fire hydrants, then use a high-fidelity simulator to determine whether the scenario is able to cause the AV to violate basic safety properties and traffic laws. Our approach, called $\mathtt{TrashFuzz}$, does not require any knowledge of how the AV and its perception system works; rather, it uses a greedy search algorithm to optimise the placement of objects and come ‘closer’ to violating signal temporal logic (STL) formulas expressing traffic laws. To ensure the realism of the generated scenarios, our algorithm is required to satisfy the constraints of various regulatory documents regarding the placement of roadside objects, including the ‘Code of Practice for Works on Public Streets’ [5, 6], the ‘Greenery Provision and Tree Conservation’ guidelines [7], as well as guidelines published by three prominent waste recycling companies, HDS, KIMBLE, and GRANGER [8, 9, 10]. Collectively, these guidelines establish a standard for the (otherwise subjective) ‘natural’ placement of everyday roadside objects such as fixed position boxes, trees, and trash bins. Figure 1 summarises the workflow of $\mathtt{TrashFuzz}$.
We implemented $\mathtt{TrashFuzz}$ for a state-of-the-art ADS, Baidu Apollo (version 7.0 [18]), and the high-fidelity Unity-based LGSVL simulator [19]. We utilise the Unity Asset Store [20] to generate a diversity of everyday roadside objects within LGSVL, including trash bins in various shapes and colours, benches constructed with different materials, as well as fire hydrants, trees, and trash bags. $\mathtt{TrashFuzz}$ ensures these objects are placed according to rules extracted from the previously mentioned regulatory guidelines, then searches for scenarios in which the objects’ placement can trigger the AV to violate Signal Temporal Logic (STL)-based specifications of traffic laws. To evaluate the approach, we conducted a comprehensive comparative analysis against of our fuzzing algorithm and found that $\mathtt{TrashFuzz}$ is efficient in generating scenarios that lead to various violations of the specifications. In addition, we found that 15 of the 24 traffic laws were violated by Apollo after solely manipulating the natural positions of roadside objects. Most of these attack scenarios induced hesitation in Apollo, causing it to fail to proceed. A few scenarios compelled Apollo to adopt a more aggressive driving behaviour, thereby significantly increasing the risk of potentially catastrophic accidents. This includes a scenario in which an otherwise benign placement of bins overloads Apollo’s perception system, causing it to perceive a red traffic light as green and illegally drive through it.
The overall results of this paper suggest that regulatory guidelines for the placement of objects on streets may need to consider the presence of AVs on the road, and that testing the resilience of AVs must also take into account anomalies arising from the placement of roadside objects.
The structure of this paper is organized as follows. In Section 2, we provide the necessary background information and define the attack goal, along with an introduction to our threat model. Section 3, we present our fuzzing algorithm. We present the evaluation of our algorithm in Section 4. Lastly, we review related work in Section 5 and conclude in Section 6.
2 Background and Problem
@@
2.1 ADS Perception Systems
The perception systems of state-of-art ADSs such as Apollo [21], Autoware [22], and Waymo [2] have a number of common characteristics. These include the integration of multiple sensors and the utilisation of Deep Neural Networks (DNNs) for classification and segmentation tasks.
Integration of Multiple Sensors. The perception module assumes a pivotal role in ADSs. These advanced systems incorporate a multitude of sensors to improve the accuracy and resilience of perception. To the best of our knowledge, current Level-4 AVs, exemplified by Apollo [21] and Waymo [2], mandate a minimum of two distinct input sources for perception. This amalgamation of sensors, cameras, LiDAR, and radar, ensures ADSs are able to gather comprehensive and reliable information about their surrounding environments.
While this multi-sensor approach significantly improves an ADS’ perception ability, it does not guarantee its resilience in all circumstances. One example is MSF-ADV [11], which demonstrates the ability to create invisible adversarial objects that deceive both the camera and LiDAR detection algorithm.
DNN-based classification/segmentation. DNNs have a distinct advantage at processing data that can be represented as vectors. They are effective at learning complex patterns and relationships within high-dimensional data, and thus can play an important role in many tasks related to ADS perception, including object detection based on both camera and LiDAR data points. Apollo [21] and Autoware [22] are examples of Level-4 ADSs that utilise DNNs to process this data.
It is well-known, however, that DNNs may be susceptible to adversarial attacks in which inputs are manipulated to trigger a misperception. For example, an adversarial object placed in the road’s centre may be ignored by DNN-based camera and LiDAR object detection algorithms [11]; the strategic placement of adversarial patches on a roadside billboard may have a notable impact on the driving behaviour of DNN-based algorithms used by the ADS [23, 24, 25]; a dirty patch on the road surface can cause DNN-based Automated Lane Centering to deviate from the intended steering angle [14]; and the presence of an adversarial patch on the vehicle in front can result in a malfunction of the DNN-based depth estimation algorithm within ADS [15]. Note that these adversarial attacks primarily fall under the category of white-box attacks (i.e. involving patches/objects with suspicious shapes).
However, ensuring the resilience of DNNs against these attacks is a difficult problem. While a few methods [26, 27, 28, 29] have demonstrated effectiveness against specific categories of adversarial attacks, none of them can serve as a universal solution for countering all types of attacks as highlighted in surveys on adversarial attacks and defences strategies [30, 31], meaning that it remains important to systematically test ADSs against them.
2.2 Attack Goal and Threat Model
Attack Goal. We propose a black box adversarial attack on the perception systems of ADSs, with the goal of inducing behaviour in which the AV violates traffic laws. This must be achieved by objects that follow conventional road design guidelines and appear benign to humans (i.e. no patches or unusual shapes). Such scenarios are termed ‘natural’: they are subsets of all possible scenarios and are characterised by restrictions on the placement of objects like bins, billboards, and greenery. We formulate the attack goal as follows:
@@ 2.2 Attack Goal and Threat Model
Threat Model. Our attack assumes a black box setting, in which the attacker has no access to the internal details of the perception system and its models. The attacker is only able to observe the outputs of the perception system and chassis information of the ADS. Furthermore, the attacker does not possess any knowledge of the inner architecture of the ADS or any algorithms that it utilises. This ensures that the attack can be used to test the resilience of any ADS by treating its perception system as a black box. The attacker’s ability is limited to modifying and moving roadside objects. They are further restricted to ‘natural’ object placements that satisfy government regulations [32] and company guidelines [8, 9, 10] ensuring scenarios are realistic. Furthermore, the attacker cannot utilise adversarial patches, as these appear as visually unnatural and require whitebox knowledge of the perception models, which in our case the attacker does not have.
2.3 Specifying Safety Properties
@@
Our attack goal is to cause misperceptions that result in the violation of safety properties. In the context of AVs, safety should not simply mean the absence of collisions, but also adherence to the rules of the road that drivers are supposed to abide by. To that end, we adopt the property specification language used by LawBreaker [17], as well as the project’s existing specifications of the traffic laws of China and Singapore. The specification language is based on Signal Temporal Logic (STL), and is evaluated with respect to traces of scenes, providing a way to automatically determine whether a tester-defined property was violated or not in a simulated run of the ADS. We highlight the key features of the specification language below (the full syntax and semantics is given in [17]).
ϕ:=assignitalic-$\phi\;:=$ $\;\mu\;|\;\neg\phi\;|\;\phi_{1}\lor\phi_{2}\;|\;\phi_{1}\land\phi_{2}\;|\;\phi_{1}\;\mathtt{U_{I}}\;\phi_{2}\;$
$\mu\;:=$ $\;f(x_{0},x_{1},\cdots,x_{k})\sim 0\ \ \sim\;:=>\;|\geq\;|<\;|\leq\;|\neq\;|=;$
Figure 2: Specification language syntax, where $\phi$, $\phi_{1}$ and $\phi_{2}$ are STL formulas, $I$ is an interval, and $f$ is a multivariate linear continuous function over language variables $x_{i}$
The high-level syntax of the language is shown in Figure 2. A time interval $I$ is of the form $[l,u]$, where $l$ and $u$ are respectively the lower and upper bounds of the interval. Following convention, we write $\Diamond_{I}\;\phi$ to denote $true\;\mathtt{U}_{I}\;\phi$; and $\Box_{I}\;\phi$ to denote $\neg\;\Diamond_{I}\;\neg\phi$. Intuitively, $\mathtt{U}$, $\Box$, and $\Diamond$ are modal operators that are respectively interpreted as ‘until’, ‘always’, and ‘eventually’. We omit the time interval when it is $[0,\infty]$.
In general, $\mu$ can be regarded as a proposition of the form $f(x_{0},x_{1},\cdots,x_{k})\sim 0$, where $f$ is a multivariate function and $x_{i}$ for all $i$ in $[0,k]$ is a variable supported in the language.
Example 2.1.
@@
Suppose we have a signal variable $speed=\langle speed(0),speed(1),\dots,speed(n)\rangle$, which represents the autonomous vehicle’s speed throughout its journey. Then, we can create a simple Boolean Expression $\mu=speed(t)<100$ to check whether the speed of the vehicle is larger than $100km/h$, as illustrated in Fig 2. Note that $\mu$ can be regarded as a proposition of the form $0100-speed(t)>$ 0 or $speed(t)-100<0$. To verify whether $\mu$ holds true at all time steps, we can straightforwardly incorporate the temporal logic symbol “always”, resulting in the formulation of $\varphi=\Box(speed<100)$.
@@
A specification is evaluated with respect to a trace $\pi$ of scenes, denoted as $\pi=\langle\pi_{0},\pi_{1},\pi_{2}\ldots,\pi_{n}\rangle$, where each scene $\pi_{i}$ is a valuation of the propositions at time step $i$ and $\pi_{0}$ reflects the state at the start of a simulation. The language follows the standard semantics of STL (see e.g. [33]).
2.4 Design Challenges
@@
C1. Restricting to ‘natural’ scenarios. A key objective of our work is to develop attack scenarios that appear as completely ‘natural’ to humans. Previous work, such as NADE [34], has focused on replicating natural driving behaviours of vehicles. In contrast, our focus is on the strategic placement of typical roadside objects, including (but not limited to) trash bins, trees, and fire hydrants. A significant challenge is knowing when the placement of such objects crosses the boundary between something innocuous to something suspicious. To that end, we constrain the placement of objects according to existing regulations and guidelines (e.g. [32, 8, 9, 10]), which we use to standardise the notion of natural scenarios.
In addition, by restricting our scenarios to natural ones, we impose constraints that significantly reduce the available space of modifications, requiring testing algorithms that navigate them more creatively. We cannot arbitrarily modify the environment that the ADS is driving in. For example, placing a trash in directly in the middle of the road should be forbidden in our testing context, as this is clearly an unnatural scenario that violates standard guidelines on bin placement.
C2. Large black-box search space. Despite restricting the placement of objects to conventional road design guidelines, there remains an enormous number of possible combinations of placements, making the search space vast and challenging to navigate, particularly in a black box setting. For example, even when limiting to only 10 permissible placement locations, if each of them can be occupied by one of 10 different possible objects, we already have $10^{10}$ combinations of object placements. Hence, an effective search algorithm is necessary.
3 Attack Design
We propose $\mathtt{TrashFuzz}$ to address the challenges highlighted in Section 2.4 as described in the following.
3.1 Generating Natural Scenarios (C1)
TABLE I: Specifications for ‘natural’ scenarios, based on public road design guidelines

### v3

# Natural Adversaries: Fuzzing Autonomous Vehicles with Realistic Roadside Object Placements

@@ Abstract
The emergence of Autonomous Vehicles (AVs) has spurred research into testing the resilience of their perception systems, i.e., ensuring that they are not susceptible to critical misjudgements. It is important that these systems are tested not only with respect to other vehicles on the road, but also with respect to objects placed on the roadside. Trash bins, billboards, and greenery are examples of such objects, typically positioned according to guidelines developed for the human visual system, which may not align perfectly with the needs of AVs. Existing tests, however, usually focus on adversarial objects with conspicuous shapes or patches, which are ultimately unrealistic due to their unnatural appearance and reliance on white-box knowledge. In this work, we introduce a black-box attack on AV perception systems that creates realistic adversarial scenarios (i.e., satisfying road design guidelines) by manipulating the positions of common roadside objects and without resorting to “unnatural” adversarial patches. In particular, we propose TrashFuzz, a fuzzing algorithm that finds scenarios in which the placement of these objects leads to substantial AV misperceptions—such as mistaking a traffic light’s colour—with the overall goal of causing traffic-law violations. To ensure realism, these scenarios must satisfy several rules encoding regulatory guidelines governing the placement of objects on public streets. We implemented and evaluated these attacks on the Apollo autonomous driving system, finding that TrashFuzz induced violations of 15 out of 24 traffic laws.
@@
Autonomous vehicles; perception systems; fuzz testing; adversarial scenarios; safety-critical systems.
I Introduction
Autonomous Vehicles (AVs) are rapidly advancing, with several Level-4 systems successfully deployed in real traffic environments [45, 59, 3, 54]. This progress has intensified interest in evaluating the resilience of their perception systems, which must reliably interpret not only vehicles and road obstacles but also roadside objects. Everyday items such as trash bins, billboards, and greenery are placed according to guidelines [29, 48, 36, 22, 26, 16] designed for human drivers, and may not align with the requirements of AV perception.
Most existing adversarial tests focus on conspicuous patches or irregularly shaped objects [5, 17, 67, 47, 10, 35], such as MSF-ADV’s visibly altered traffic cone [5]. These attacks typically pursue objectives such as making an object “ignored” or mislocalised and often require white-box access to perception models. While such misperceptions may expose safety issues, their unrealistic appearance and impractical assumptions limit applicability on real roads.
These limitations motivate two questions. First, can AV perception be deceived by natural, guideline-compliant roadside objects—those appearing entirely benign to human observers? Second, if such misperceptions occur, can they escalate into serious safety issues, such as collisions or violations of traffic laws [50]? Demonstrating these rare but plausible scenarios in simulation is essential before AVs encounter them in the wild.
@@ I Introduction
Figure 1: TrashFuzz workflow: mutating roadside-object placements to find natural scenarios that induce an autonomous vehicle to violate traffic laws
To address these questions, we introduce TrashFuzz, a black-box fuzzing approach that systematically searches for guideline-compliant object placements capable of misleading an Autonomous Driving System (ADS) into unsafe behaviour. TrashFuzz generates scenarios using naturally shaped, patch-free objects (e.g., bins, benches, trees, hydrants), evaluates them in a high-fidelity simulator, and uses a greedy search to move “closer” to violating Signal Temporal Logic (STL) specifications of traffic laws. The fuzzer is constrained by regulatory documents—including the Code of Practice for Works on Public Streets [29, 48], greenery-provision guidelines [36], and placement rules from major waste-management companies [22, 26, 16]—ensuring all scenarios remain realistic. Figure 1 illustrates the workflow.
Unlike LawBreaker [50], which mutates the behaviour of agents (e.g., other vehicles or pedestrians) to induce traffic-law violations, TrashFuzz treats regulation-compliant roadside object layouts themselves as the adversarial input space. It introduces naturalness as a constraint, formalising real-world placement guidelines as executable rules that are enforced during scenario generation. Within this constrained space, TrashFuzz performs gradient-guided black-box optimisation over object types, positions, and orientations to minimise STL robustness and induce perception-driven law violations.
We implement TrashFuzz for the Baidu Apollo ADS [3] with the Unity-based LGSVL simulator [44], using the Unity Asset Store [56] to provide a wide range of everyday roadside objects. TrashFuzz automatically enforces placement constraints and searches for scenarios that trigger STL-defined law violations. In our evaluation, TrashFuzz efficiently generated scenarios that induced Apollo to violate 15 of 24 testable traffic laws. Most cases caused hesitation or failure to proceed; others induced aggressive behaviour, including one scenario in which an innocuous arrangement of bins overloaded the perception system and caused the ADS to misinterpret a red light as green.
These findings highlight that existing placement guidelines, designed primarily for human drivers, may be inadequate in the presence of AVs, and that testing pipelines must account for subtler, object-induced anomalies.
II Background and Problem
@@
II-A ADS Perception Systems
State-of-the-art ADSs such as Apollo [4], Autoware [2], and Waymo [59] share two key perception characteristics: multi-sensor integration and the use of Deep Neural Networks (DNNs) for classification and segmentation.
Integration of Multiple Sensors. Modern Level-4 AVs typically fuse data from at least two sensor modalities—commonly cameras, LiDAR, and radar—to obtain a more accurate and robust view of the environment. While this multi-sensor setup enhances perception reliability, it does not guarantee resilience. For example, MSF-ADV [5] shows that carefully crafted objects can deceive both camera- and LiDAR-based detectors despite the presence of multiple sensors.
DNN-based Classification and Segmentation. DNNs are widely used in ADS perception pipelines for detecting and segmenting objects from camera and LiDAR data [4, 2]. However, DNNs are known to be vulnerable to adversarial manipulation. Prior work demonstrates that adversarial objects or patches can cause detectors to ignore obstacles [5], misinterpret roadside billboards [63, 42, 43], deviate lane-centering controllers [47], or corrupt depth estimation [10]. These examples are largely white-box attacks involving conspicuous or irregular adversarial patterns.
Defending DNNs against such attacks remains challenging. Although several defence mechanisms [28, 21, 41, 46] show partial effectiveness, surveys [7, 8] emphasise that no universal solution exists. Thus, systematic testing of ADS perception systems continues to be essential.
II-B Attack Goal and Threat Model
Attack Goal. We propose a black-box adversarial attack on the perception systems of ADSs, with the goal of inducing behaviour in which the AV violates traffic laws. This must be achieved by objects that follow conventional road design guidelines and appear benign to humans (i.e., no patches or unusual shapes). Such scenarios are termed “natural”: they are subsets of all possible scenarios and are characterised by restrictions on the placement of objects like bins, billboards, and greenery. We formulate the attack goal as follows:
@@ II-B Attack Goal and Threat Model
Threat Model. Our attack assumes a black-box setting, in which the attacker has no access to the internal details of the perception system and its models. The attacker is only able to observe the outputs of the perception system and chassis information of the ADS. Furthermore, the attacker does not possess any knowledge of the inner architecture of the ADS or any algorithms that it utilises. This ensures that the attack can be used to test the resilience of any ADS by treating its perception system as a black box. The attacker’s ability is limited to modifying and moving roadside objects. They are further restricted to natural object placements that satisfy government regulations [49] and company guidelines [22, 26, 16] ensuring scenarios are realistic. Furthermore, the attacker cannot utilise adversarial patches, as these appear as visually unnatural and require white-box knowledge of the perception models, which in our case the attacker does not have.
II-C Specifying Safety Properties
@@
Our attack goal is to cause misperceptions that result in the violation of safety properties. In the context of AVs, safety should not simply mean the absence of collisions, but also adherence to the rules of the road that drivers are supposed to abide by. To this end, we adopt the property specification language used by LawBreaker [50], as well as the project’s existing specifications of the traffic laws of China and Singapore. The specification language is based on STL formulas, and is evaluated with respect to traces of scenes, providing a way to automatically determine whether a tester-defined property was violated or not in a simulated run of the ADS. We highlight the key features of the specification language below (the full syntax and semantics is given in [50]).
$\phi\;::=$ $\;\mu\;|\;\neg\phi\;|\;\phi_{1}\lor\phi_{2}\;|\;\phi_{1}\land\phi_{2}\;|\;\phi_{1}\;\mathcal{U}_{I}\;\phi_{2}$
$\mu\;::=$ $\;f(x_{0},x_{1},\ldots,x_{k})\;\sim\;0,\qquad\sim\in\{>,\geq,<,\leq,=,\neq\}$
Figure 2: Syntax of our specification language, a fragment of Signal Temporal Logic (STL). Here, $\phi$ denotes a formula, $\mu$ is an atomic predicate over real-valued functions $f$ of variables $x_{i}$, and $I=[l,u]$ is a bounded or unbounded time interval
The high-level syntax of our specification language is formalised in Figure 2. Temporal formulas $\phi$ are built from atomic predicates $\mu$ using Boolean connectives and the bounded temporal operator $\mathcal{U}_{I}$ (interpreted as ‘until over interval $I$’). A time interval $I$ is of the form $[l,u]$, where $l$ and $u$ are non-negative real-valued bounds. We adopt standard abbreviations: the eventually operator $\Diamond_{I}\phi$ is syntactic sugar for $true\;\mathcal{U}_{I}\;\phi$, and the always operator $\Box_{I}\phi$ abbreviates $\neg\Diamond_{I}\neg\phi$. When $I=[0,\infty]$, the interval may be omitted.
An atomic predicate $\mu$ is a relational constraint over a multivariate function $f(x_{0},x_{1},\ldots,x_{k})$ and a constant threshold 0, using standard relational operators. These variables $x_{i}$ are language-defined signals or features, such as vehicle speed, object distance, or perception confidence scores.
Example II.1.
@@
Suppose we have a signal variable $speed=\langle speed(0),speed(1),\dots,speed(n)\rangle$, which represents the autonomous vehicle’s speed throughout its journey. Then, we can create a simple Boolean expression $\mu=speed(t)<100$ to check whether the speed of the vehicle is larger than $100km/h$, as illustrated in Fig 2. Note that $\mu$ can be regarded as a proposition of the form 100speed(t)>0100-speed(t)>0 or speed(t)100<0speed(t)-100<0. To verify whether $\mu$ holds true at all time steps, we can straightforwardly incorporate the temporal logic symbol “always”, resulting in the formulation of $\varphi=\Box(speed<100)$.
@@
A specification is evaluated with respect to a trace $\pi$ of scenes, denoted as $\pi=\langle\pi_{0},\pi_{1},\pi_{2}\ldots,\pi_{n}\rangle$, where each scene $\pi_{i}$ is a valuation of the propositions at time step $i$ and $\pi_{0}$ reflects the state at the start of a simulation. The language follows the standard semantics of STL (see e.g., [33]).
II-D Design Challenges
@@
C1. Restricting to natural scenarios. A key objective of our work is to develop attack scenarios that appear as completely natural to humans. Previous work, such as NADE [15], has focused on replicating natural driving behaviours of vehicles. In contrast, our focus is on the strategic placement of typical roadside objects, including (but not limited to) trash bins, trees, and fire hydrants. A significant challenge is knowing when the placement of such objects crosses the boundary between something innocuous to something suspicious. To that end, we constrain the placement of objects according to existing regulations and guidelines (e.g., [49, 22, 26, 16]), which we use to standardise the notion of natural scenarios.
In addition, by restricting our scenarios to natural ones, we impose constraints that significantly reduce the available space of modifications, requiring testing algorithms that navigate them more creatively. We cannot arbitrarily modify the environment that the ADS is driving in. For example, placing a trash bin directly in the middle of the road should be forbidden in our testing context, as this is clearly an unnatural scenario that violates standard guidelines on bin placement.
C2. Large black-box search space. Despite restricting the placement of objects to conventional road design guidelines, there remains an enormous number of possible combinations of placements, making the search space vast and challenging to navigate, particularly in a black-box setting. For example, even when limiting to only 10 permissible placement locations, if each of them can be occupied by one of 10 different possible objects, we already have $10^{10}$ combinations of object placements. Hence, an effective search algorithm is necessary.
III TrashFuzz Methodology
We propose TrashFuzz to address these challenges. Figure 3 provides an overview, which we elaborate on below.
Figure 3: Overview of the TrashFuzz approach
III-A Generating Natural Scenarios (C1)
TABLE I: Specifications for natural scenarios, based on public road design guidelines